OpenStudy (anonymous):

find the exponential function f(x)=ab^x, if f(0) = 64, and f(2)=4

OpenStudy (anonymous):

f(0) = 64. At zero, function is like a . b ^0 We know that anything to the power 0 = 1 therefore we get a = 64. Now at 2, function is like, a . b ^ 2 = 4 a is 64. Therefore, b ^2 = 1/16 = 4 ^ (-2) Comparing it gives us b = 4. There fore function is, 64 . 4^x